A total of $15,000 is invested in two corporate bonds that pay 5% and 7% simple annual interest. The investor wants to earn $880 in interest per year from the bonds. How much should be invested in each bond?
Answer by mananth(16949)

5% --------x
7% --------15000-x
...
Interest = 880
5x+7(15000-x)=880*100
5x+105000-7x=88000
-2x=-17000
/-2
x=8500 at 5%
balance at7%